Suppose AB is the diameter of a circle, where A =(-1,1) and B = (-4,-5). What is the equation of the circle?
1
Expert's answer
2019-06-11T09:36:29-0400

The center-radius form of the circle equation is in the format (x-a)2+(x-b)2=r2, with the center being at the point O(a, b) and the radius being "r ". If AB is the diameter of a circle, where A =(-1,1) and B = (-4,-5), then the center O(a, b) is a midpoint of AB.

a=(-1+(-4))/2=-2.5, b=(1+(-5))/2=-2, then the center is O(-2.5,-2). 

r=AB/2

r2=AB2/4

AB2= (-4-(-1))2+(-5-1)2=(-5)2+(-6)=25+36=61

 r2=61/4=15.25

r2=15.25

(x-(-2.5))2+(y-(-2))2=15.25

The equation of the circle is (x+2.5)2+(y+2)2=15.25